# Fast Matrix Multiplications for Lookup Table-Quantized LLMs

### v3

@@ 4.2 End-to-End LLM Benchmarks
As an application of FLUTE, we experiment with quantizing LLaMA3-8B and LLaMA3-70B. The LLaMA3 family of models has been found to be more difficult to quantize than other open source models Huang et al. (2024), and thus presents a testing ground for different quantization strategies.
@@
Table 2: Evaluation of post-training quantization on LLaMA3-8B and LLaMA3-70B. The RTN, GPTQ Frantar et al. (2022), AWQ Lin et al. (2023) results are from Huang et al. (2024); the rest are from our implementations. All non-NF methods use uniform weight quantization.
@@ Appendix A Appendix
Table 3: Detailed evaluation of post-training quantization on LLaMA3-8B. The RTN, GPTQ Frantar et al. (2022), AWQ Lin et al. (2023) results are from Huang et al. (2024); the rest are from our implementations. All non-NormalFloat methods use uniform weight quantization.
@@
Table 4: Detailed evaluation of post-training quantization on LLaMA3-70B. The RTN, GPTQ Frantar et al. (2022), AWQ Lin et al. (2023) results are from Huang et al. (2024); the rest are from our implementations. All non-NormalFloat methods use uniform weight quantization.

### v4

@@ 4.2 End-to-End LLM Benchmarks
As an application of FLUTE, we experiment with quantizing LLaMA3-8B and LLaMA3-70B. The LLaMA3 family of models has been found to be more difficult to quantize than other open source models Chen et al. (2024), and thus presents a testing ground for different quantization strategies.
@@
Table 2: Evaluation of post-training quantization on LLaMA3-8B and LLaMA3-70B. The RTN, GPTQ Frantar et al. (2022), AWQ Lin et al. (2023) results are from Chen et al. (2024); the rest are from our implementations. All non-NF methods use uniform weight quantization.
@@ Appendix A Appendix
Table 3: Detailed evaluation of post-training quantization on LLaMA3-8B. The RTN, GPTQ Frantar et al. (2022), AWQ Lin et al. (2023) results are from Chen et al. (2024); the rest are from our implementations. All non-NormalFloat methods use uniform weight quantization.
@@
Table 4: Detailed evaluation of post-training quantization on LLaMA3-70B. The RTN, GPTQ Frantar et al. (2022), AWQ Lin et al. (2023) results are from Chen et al. (2024); the rest are from our implementations. All non-NormalFloat methods use uniform weight quantization.